Ok, I got -8 for this problem. I'm not certain if its right.
5x+3+5x+9=9x+4

Answer by Alan3354(21605)

5x+3+5x+9=9x+4
10x + 12 = 9x + 4
x + 12 = 4
x = -8
---------------